OpenStudy (anonymous):

Please help. Find all the complex zeros f(x)=x^3-10x^2+44x-69. I don't know what to do after this part below.

OpenStudy (anonymous):



OpenStudy (anonymous):

well you're doing it correct... once you got zero (0) remainder... next will be a second degree equation which is much easier to solve since we have the quadratic formula for solving the zeros of the function...

OpenStudy (anonymous):

So I just keep going until I get zero? What do I do with the remainder I have now?

OpenStudy (anonymous):

if you have a remainder, it means it is not a zero of the function, try another factor of 69...

OpenStudy (alekos):

There is no remainder. the last term on the second line should be 69 ( 23 x 3 )

OpenStudy (anonymous):

then the resulting numbers of each column would the constants of the 2nd degree or quadratic equation... if 69 is a factor... the resulting factor of the above equation will be... (x - 69)(ax^2 + bx + c)... where a, b, and c are the result of synthetic divisions you made above...

OpenStudy (anonymous):

or is it 3 not 69?

OpenStudy (anonymous):

@alekos Oh, I caught that mistake a while ago. But I'm unsure of what I should do after getting x^2-7x+23. How would I find the real zero and complex zeros? Do I use the quadratic formulas?

hero (hero):

Yes, just simply use the quadratic formula. Remeber \(\sqrt{-1} = i\)

OpenStudy (anonymous):

Okay thanks @Hero

OpenStudy (anonymous):

I'm sorry, What's the \[\sqrt{-43}\] ? lol

hero (hero):

Recall this square root property: \(\sqrt{ab} = \sqrt{a}\sqrt{b}\) In this case, \(a = -1\) and \(b = 43\) and \(\sqrt{-43} = \sqrt{-1}\sqrt{43}\)

OpenStudy (anonymous):

Oh I did that right then. I thought I was wrong, lol. After I've done that, I'm left with \[\frac{ 7\pm \sqrt{-1}\sqrt{43} }{ 2}\] How would that help me find zeros?

hero (hero):

As I said before \(\sqrt{-1} = i\)

hero (hero):

You use the quadratic formula to find the zeroes of a quadratic function. Once you have simplied it completely, the result represents the zeroes of the function.

OpenStudy (anonymous):

I never really understood imaginary #s so I don't understand how I'm supposed to simplify. \[\frac{ 7\pm \sqrt{43i} }{ 2 }\]

hero (hero):

Actually, you should have written it as \(\dfrac{ 7\pm \sqrt{43}i }{ 2 }\) or \(\dfrac{ 7\pm i\sqrt{43} }{ 2 }\) The \(i\) doesn't go under the square root.

OpenStudy (anonymous):

Oh. Thanks for helping me. :) Really appreciate it. My goal was to find all the complex zeros, but I still have no idea what the zeros are after doing all this work, lol. This is a difficult problem for me. I'm sorry, lol.

OpenStudy (anonymous):

additional... the zeros you are referring to are the roots of the functions, the first root you got is a real number, and the last two are of complex numbers having an "i" term known as imaginary numbers...

hero (hero):

I meant to say earlier that the quadratic formula would be used to find the complex zeroes of the quadratic function.

hero (hero):

You found them

OpenStudy (anonymous):

So the real number is 3 and the complex is \[\frac{ 7\pm i \sqrt{43} }{ 2 }\] Am I right?

OpenStudy (anonymous):

Alright! Thank you so much @Hero and @Orion1213 for helping me! You guys are awesome!

hero (hero):

Actually, I only covered with you the complex part of it.

hero (hero):

There's a way to figure out if 3 is a zero.

OpenStudy (anonymous):

Wasn't that done with the synthetic division though?

hero (hero):

Your original function is f(x)=x^3-10x^2+44x-69 If f(3) = 0, then 3 is a zero. If f(-3) = 0, then -3 is a zero. It's one of the two.

hero (hero):

Also, you can just graph the function to figure it out quickly

OpenStudy (anonymous):

How do you graph it? I don't have a graphing calculator on me.

OpenStudy (anonymous):

Oh you've already written the problem. That's the graph?

OpenStudy (anonymous):

Now everything is finished right?

hero (hero):

Yes, you found all the complex zeroes. By the way, 3 is also a complex zero. 3 + 0i So include 3 in your solution of complex numbers.

OpenStudy (anonymous):

So when I write my answers down it will look like this? Complex: \[\pm3\] and \[\frac{ 7pmi \sqrt{43} }{ 2 }\]

OpenStudy (anonymous):

woops messed up on writing the last one. You get it, lol.

hero (hero):

-3 is not a zero.

OpenStudy (anonymous):

Oh thanks.